OpenStudy (anonymous):

a triangle has side lenghts of 15, 20, and 25. what type of triangle is it?

OpenStudy (anonymous):

right angled

OpenStudy (anonymous):

explain

OpenStudy (angela210793):

If 25^2=15^2+20^2 then it is a right angled triangle

OpenStudy (anonymous):

because 25^2 = 20^2 + 15^2 (Pythagoras)

OpenStudy (anonymous):

the 2 lowest right angled triangles with whole numbers are 3,4,5 and 5,12,13 15,20,25 is a multiple of the first one